Mr. JOnes was charged the following rates for his phone bill:
$1.85 for 37 minutes
$3.75 for 1 hour 15 minutes (75 minutes)
a) Put this information into (x,y) pairs. Which one is the independent variable?
b) What is the rate charged per minute?
c) IS there a fee for having the phone?_____ If so, what is it?
d) What is the equation that can be used to find the price for any amount of minutes used?
e) What would be the cost if Mr. Jones used 118 minutes?

Answer by stanbon(57203)

Mr. JOnes was charged the following rates for his phone bill:
$1.85 for 37 minutes
$3.75 for 1 hour 15 minutes (75 minutes)
-----------
You have two points: (37,1.85), (75,3.75)
a) Put this information into (x,y) pairs. Which one is the independent variable?
The independent variable is time of use.
b) What is the rate charged per minute?
Find the equation relating minutes to cost, as follows:
slope = [3.75-1.85]/[75-37]=1.90/38=0.05
0.05 is the rate charged per minute
---------------
intercept: 3.75=0.05(75)+b ; 3.75-0.05(75)+b; b=0
EQUATION: Cost=0.05(time in minutes)+0
--------
c) IS there a fee for having the phone?_____ If so, what is it?
Cost fo zero minutes = 0.05(0)=0
So there is no cost for just having the phone.
d) What is the equation that can be used to find the price for any amount of minutes used?
Cost=0.05(minutes used)
e) What would be the cost if Mr. Jones used 118 minutes?
Cost=0.05(118)=$5.90
Cheers,
Stan H.